OpenStudy (anonymous):

What is the product of 3/5 x 4/9 written in simplest form? * A. 4/5 B. 4/15 C. 1/2 D. 7/15

Mehek (mehek14):

3*4 = 12 5 * 9 = 45 12 and 45 have the gcf 3 12/3 = 4 45/3 = 15 4/15 is in simplest form

OpenStudy (anonymous):

Yes the answer is B

OpenStudy (anonymous):

thanks